OpenStudy (anonymous):

If x=2 and r= 5, then y= what? If r=25 and y= 15, then x= what? If x=10 and y= 6, then whats r?

OpenStudy (lgbasallote):

\[r = \sqrt{x^2 + y^2}\] \[r = \sqrt{8^2 + 15^2}\]

OpenStudy (lgbasallote):

got that?